# A Non-Traditional Approach to Assisting Data Address Translation

### v1

@@ VI Concluding Remarks
We presented a novel technique that uses the PC of a load instruction, rather than the address of the referenced data, to obtain an PTE for the referenced data. The key observation driving the use of a PC rather than the data address is that often a dynamic instance of a load instruction accesses the same virtual page as the prior dynamic instance, thus needing the same PTE. This technique can only be profitably used by a small subset of the static loads in the program. We further observed that only a small subset of the static loads accounts for most of the address translation misses with conventional DTLBs. Combining these two observations we proposed PC-Indexed Data Address Translation (PCAX), to assist conventional address translation structures. PCAX employs tables that require additional storage, which is modest given the transistor budgets in contemporary high-performance chips, but is able to achieve a significant reduction (2-3X or more) in the effective DTLB miss rate, and improve performance, while also reducing the energy for data address translation in most of the considered benchmarks.
A plethora of approaches have been proposed to reduce the DTLB miss rate, with varying degrees of effectiveness and with varying consequences for system software over the past several decades. PCAX is a purely microarchitectural solution that is highly effective, with no demands on the ISA, software and/or systems.
To limit the number of loads for which PCAX is applied, and keep the associated tables (reasonably) small, we limited the application of PCAX to loads that are likely to miss in a DTLB. It can also be applied to a few “high value” loads that would normally hit in a DTLB and thus further reduce the overall latency and energy for data address translation. Using the PC of an instruction to obtain a translation for the data that it will access also opens up the possibility of many other innovations for data address translation. We leave these to future work.

### v2

@@ VI Concluding Remarks
We presented a novel technique that uses the PC of a load instruction, rather than the address of the referenced data, to obtain an PTE for the referenced data. The key observation driving the use of a PC rather than the data address is that often a dynamic instance of a load instruction accesses the same virtual page as the prior dynamic instance, thus needing the same PTE. This technique can only be profitably used by a small subset of the static loads in the program. We further observed that only a small subset of the static loads accounts for most of the address translation misses with conventional DTLBs. Combining these two observations we proposed PC-Indexed Data Address Translation (PCAX), to assist conventional address translation structures. PCAX employs tables that require additional storage, which is modest given the transistor budgets in contemporary high-performance chips, but is able to achieve a significant reduction (2-3X or more) in the effective DTLB miss rate, and improve performance, while also reducing the energy for data address translation in most of the considered benchmarks. PCAX is a purely microarchitectural solution that is highly effective, with no demands on the ISA, software and/or systems, unlike other proposed techniques which have varying consequences for system software.
Future work includes assessing if the observations that underlie PCAX can be applied to create new prefetching or replacement schemes for an STLB, and perhaps even a CDTLB, with or without PCAX, to further improve translation efficiency.